Find dy/dx by implicit differentiation. cot(y) = 5x – 6y dy/dx = Need Help? Read It Talk to a Tutor
Answer #1

a Ans an given that Cot(y) = 5x by for implicit differentiation. en (coty) = ( 53-64) - Cosecydy = 5-6 dy –5= (cosec?y-6) dyif you happy this answer then plz give me positive rating